# Fast offline decoding with local message-passing automata

### v1

@@ 6.3 Performance guarantees for Lindbladian decoding
[…]
@@
[…]
@@
[…]
@@ 7.2 Two dimensions
We now repeat the above analysis for the 2D toric code. The left panel of Fig. 11 plots $p_{\sf log}$ as a function of $p$ for different system sizes. Since our simulations are done on the torus, where logical errors occur when anyon strings wrap around either one or both of the two nontrivial cycles, $p_{\sf log}\rightarrow 3/4$ in the limit $p\rightarrow 1/2$. The crossing points for different values of $L$ are consistent with a threshold of $p_{c}\approx 7.3\%$, which unlike the 1D case is (a few percent) smaller than the optimum value of $\approx 11\%$ set by the phase transition of the 2D random bond Ising model on the Nishimori line [21], but is a bit higher than the $\approx 6.1\%$ of the “constant-$c$” field-based decoder of Ref. [13]. The right panel of Fig. 11 plots $p_{\sf log}$ against $L$ and confirms the expected scaling $p_{\sf log}=e^{-\Theta(L)}$ for $p$ not too far below threshold. Finally, Fig. 12 shows $T_{\sf dec}$ as a function of $L$ for several values of $p<p_{c}$, which again display an excellent fit to the predicted scaling $T_{\sf dec}=\Theta(\log L)$ (as in 1D, introducing local spatial correlations into the noise does not appear to affect the results).
@@
We can give an extremely crude estimation of $p_{c}$ for i.i.d noise using the following simplistic argument. Let $\mathsf{P}_{\sf nearest\,\,paired}$ be the probability that a given anyon is paired with one of its nearest neighbors. Heuristically, our decoder should succeed when this probability is $>1/2$; in this case, moving each anyon towards its nearest neighbor is likely to induce a pairing leading to an error-free perfect matching. We compute $\mathsf{P}_{\sf nearest\,\,paired}$ numerically in Fig. 13, which shows that $\mathsf{P}_{\sf nearest\,\,paired}=1/2$ at $p\approx 6.5\%$, which is slightly smaller than the numerically computed value of the threshold probability. It would be interesting if a refinement of this naive argument could be used to get a sharp bound on $p_{c}$.
@@ 7.3 Uncoordinated desynchronization
For all cellular automata capable of highly structured computations known to the author—such as the game of life and many of the 1D automata defined by Wolfram—performing uncoordinated desynchronization completely destroys the automaton’s ability to perform meaningful computations. However, for Toom’s rule, which like the present decoder is designed only to remember $O(1)$ bits of information, uncoordinated desynchronization is provably innocuous, changing the value of $p_{c}$ but not the existence of a threshold [51, 52, 53]. While it is possible to construct local robust classical memories storing $O(1)$ bits for which uncoordinated desynchronization provably fails [54], it is still reasonable to conjecture that it may work for the present message-passing decoder under consideration.
@@
Figure 15: Performance of the uncoordinated asynchronous 2D toric code decoder under i.i.d noise of strength $p$. Left: $p_{\sf log}^{\sf un}$ vs $p$, consistent with a threshold around $p^{\sf un}_{c}\approx 5.2\%$. Right: $p_{\sf log}^{\sf un}$ vs $p$; the solid lines are fits to $p_{\sf log}=C(p/p^{\sf un}_{c})^{\gamma L}$ with $\gamma\approx 1/10$.
@@ 8 Acknowledgments
I thank Ehud Altman, Nicholas O’Dea, Ruihua Fan, Sarang Gopalakrishnan, Jeongwan Haah, Yaodong Li, Louis Paletta, Charles Stahl, Kazuaki Takasan, Zijian Wang, and Mike Zaletel for discussions and feedback, Shankar Balasubramanian and Margarita Davydova for helpful discussions about Gacs’ sparsity methods, synchronization, and collaboration on related work, and Larry Chen and Noah Goss for insights into hardware constraints. I am supported by a Miller research fellowship.
@@ A.1 Basic notions, and definitions of level-$k$ noise and error sets
In the following, we will always assume the existence of a finite $10<p<$ 1 for which the above holds.
@@
It follows from the above definition that any $(W/2,W/2+B)$-isolated point defines a $(W,B)$-cluster. However, a given cluster $C_{(W,B)}$ may contain anywhere between 00 and $|C_{(W,B)}|$ $(W/2,W/2+B)$-isolated points; examples are shown in Fig. 16. The partition of a (subset of a) subset of $\mathbb{Z}$ into clusters is as stated not necessarily unique in general, but it is if $B>W$:
@@
Let $\mathsf{N}\subset\mathbb{Z}$ be a noise realization, and $\{C_{(W,B)}^{(i)}\}$ be a collection of $(W,B)$-clusters in $\mathsf{N}$. If $B\geq W$, these clusters are disjoint: $C_{(W,B)}^{(i)}\cap C_{(W,B)}^{(j)}$ if $i\neq j$.
@@
Fix $i,j$, assume $B\geq W$, and suppose $x\in C_{(W,B)}^{(i)}\cap C_{(W,B)}^{(j)}$. Consider a point $y\in C_{(W,B)}^{(i)}$. Then $|x-y|<R$. This implies ${\sf dist}(C_{(W,B)}^{(j)},\{y\})<W$. Thus ${\sf dist}(C_{(W,B)}^{(j)},\mathsf{N}\setminus C_{(W,B)}^{(j)})<B$ since $B\geq W$, and hence we must have $i=j$. ∎
@@ A.2 The sparsity theorem in one dimension
$\mathsf{E}_{k}=\mathsf{N}_{k+1}\setminus\mathsf{N}_{k}.$ (144)
We similarly define the level-$k$ isolated error set $\mathcal{E}_{k}$ as $\mathcal{E}_{k}=\mathcal{N}_{k+1}\setminus\mathcal{N}_{k}$.
@@
$p_{k}=\max_{x\in\mathbb{Z}}\left(x\in\mathsf{N}_{k}\right),$ (147)
@@ A.3 General dimensions
$p_{k}=\max_{\mathbf{x}\in\mathbb{Z}^{D}}(\mathbf{x}\in\mathsf{N}_{k}).$ (164)
@@
The sparsity of errors at different values of $k$ is easy enough to verify numerically; the result of doing this is shown in Fig. 17.
@@ Appendix B Open boundary conditions
=$\mathcal{B}=\{\mathbf{r}\in\Lambda\,:\,r^{1}=0\}\cup\{\mathbf{r}\in\Lambda\,:\,r^{1}=L-1\}=\mathcal{B}_{L}\cup\mathcal{B}_{R}$ (170)
@@ B.1 Modifications to the decoder
We begin by describing how to modify the message passing and feedback rules in the presence of open boundaries. In this context, error correction is performed by allowing anyons to match either with other anyons, or with sites in $\mathcal{B}$. Since sites in $\mathcal{B}$ act as anyons for the purposes of matching, we simply modify the automaton rule $\mathcal{A}$ to source messages at all such sites, with a component of $\mathcal{B}$ with outward-facing unit normal ${\mathbf{\hat{a}}}$ sourcing messages that propagate along $-{\mathbf{\hat{a}}}$. These rules lead to anyons being attracted either to the anyon nearest to them or to the nearest site in $\mathcal{B}$, whichever is closer. This modification is formalized in the following definition:
@@
where we take $C^{(a)}_{\mathbf{r}\mp{\mathbf{\hat{a}}}}=\varnothing$ if $\mathbf{r}\mp{\mathbf{\hat{a}}}\not\in\Lambda$.
@@ B.2 Erosion with open boundaries
We first claim that if an anyon enters $\Lambda\setminus\mathsf{B}$, it moves along $+{\mathbf{\hat{x}}}$ until annihilating on a site in $\mathcal{B}_{R}$. To show this, consider the first anyon $a$ from $C$ to enter $\Lambda\setminus\mathsf{B}$ (if multiple anyons enter $\Lambda\setminus\mathsf{B}$ at the same time, we may resolve the degeneracy by picking the anyon closest to $\mathcal{B}_{R}$ [or one of them, if there are multiple closest anyons]). If $\mathbf{r}_{t}$ is the spatial coordinate of anyon $a$ at time $t$, then if $t$ is the time when $a$ first moves into $\Lambda\setminus\mathsf{B}$, we must have $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r}_{t},t))=0$ and
$||\mathbf{r}_{t}-\mathcal{B}_{R}||_{\infty}<||\mathbf{r}_{t}-\mathsf{A}({\sf PLF}(\mathbf{r}_{t},t))||_{\infty},$ (177)
where we have defined the distance from $\mathbf{r}$ to a set $S$ as $||\mathbf{r}-S||=\min_{\mathbf{r}^{\prime}\in S}||\mathbf{r}-\mathbf{r}^{\prime}||$ (if $S$ is a set of spacetime points, then the minimum is taken by discarding the time coordinates). At time $t+1$, $a$ has moved in the $+{\mathbf{\hat{x}}}$ direction, hence $||\mathbf{r}_{t+1}-\mathcal{B}_{R}||_{\infty}=||\mathbf{r}_{t}-\mathcal{B}_{R}||_{\infty}-1$. Given this, and given the modified feedback rules in Def. 4.3, the only way for
$||\mathbf{r}_{t+1}-\mathsf{A}({\sf PLF}(\mathbf{r}_{t+1},t))||_{\infty}<||\mathbf{r}_{t}-\mathsf{A}({\sf PLF}(\mathbf{r}_{t},t))||_{\infty}-1$ (178)
@@
is for there to be an anyon in ${\sf PLF}^{+x}(\mathbf{r}_{t},t)$ that moves along $-{\mathbf{\hat{x}}}$ at time $t$, which contradicts $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r}_{t},t))=0$. Therefore the distance from $a$ to $\mathcal{B}_{R}$ decreases at least as much as the distance between $a$ and any point in $\mathsf{A}({\sf PLF}(\mathbf{r}_{t},t))$. This argument holds for any time $t^{\prime}>t$, and $a$ therefore moves along $+{\mathbf{\hat{x}}}$ until annihilating on a site in $\mathcal{B}_{R}$.
Now consider an anyon $a^{\prime}$ which leaves $\mathsf{B}$ but which is not among the first anyons in $C$ to do so. That $a^{\prime}$ must also move along $+{\mathbf{\hat{x}}}$ until annihilating at $\mathcal{B}_{R}$ can be shown by induction using the result of the above paragraph as a base case. Suppose the claim is true for the first $k$ anyons to leave $\mathsf{B}$. If $a^{\prime}$ is the $(k+1)$st anyon to leave $\mathsf{B}$ (with degeneracies broken as before) and it does so at $(\mathbf{r}^{\prime}_{t^{\prime}},t^{\prime})$, then either $||\mathbf{r}^{\prime}_{t^{\prime}}-\mathsf{A}({\sf PLF}(\mathbf{r}^{\prime}_{t^{\prime}},t^{\prime}))||_{\infty}=||\mathbf{r}^{\prime}_{t^{\prime}}-\mathsf{A}({\sf PLF}^{+x}(\mathbf{r}^{\prime}_{t^{\prime}},t^{\prime}))||_{\infty}$, or $||\mathbf{r}^{\prime}_{t^{\prime}}-\mathcal{B}_{R}||_{\infty}<||\mathbf{r}^{\prime}_{t^{\prime}}-\mathsf{A}({\sf PLF}(\mathbf{r}^{\prime}_{t^{\prime}},t^{\prime}))||_{\infty}$. If the latter case holds, then the inductive hypothesis and the properties of the feedback rules in Def. 4.3 mean that
$||\mathbf{r}^{\prime}_{t^{\prime}+1}-\mathsf{A}({\sf PLF}^{+x}(\mathbf{r}^{\prime}_{t^{\prime}+1},t^{\prime}+1))||_{\infty}=||\mathbf{r}^{\prime}_{t^{\prime}+1}-\mathsf{A}({\sf PLF}^{+x}(\mathbf{r}^{\prime}_{t^{\prime}+1},t^{\prime}+1))||_{\infty}$ (179)
provided $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r}^{\prime}_{t^{\prime}+1},t^{\prime}+1))=1$ (and if not, then we must have $||\mathbf{r}^{\prime}_{t^{\prime}+1}-\mathcal{B}_{R}||_{\infty}<||\mathbf{r}^{\prime}_{t^{\prime}+1}-{\sf PLF}(\mathbf{r}^{\prime}_{t^{\prime}+1},t^{\prime}+1)||_{\infty}$, which reduces to the latter case considered above), and that
$||\mathbf{r}^{\prime}_{t^{\prime}}-\mathsf{A}({\sf PLF}^{+x}(\mathbf{r}^{\prime}_{t^{\prime}},t^{\prime}))||_{\infty}\leq||\mathbf{r}^{\prime}_{t^{\prime}+1}-\mathsf{A}({\sf PLF}(\mathbf{r}^{\prime}_{t^{\prime}+1},t^{\prime}+1))||_{\infty}.$ (180)
@@
Therefore $a^{\prime}$ must continue to move along ${\mathbf{\hat{x}}}$ at the next time step; since this argument holds for all $t^{\prime}$, it shows the claim.
@@
To give a (loose) upper bound on the erosion time, we use the same strategy as in Lemma 4.4. Consider anyons at the slice $\mathsf{S}_{k}=\mathsf{B}\cap\{\mathbf{r}\,:\,r^{1}=L-1-x+W/2-k\}$ (so that $\mathsf{S}_{0}$ is the single point of $\mathsf{B}$ closest to $\mathcal{B}_{R}$, and $\mathsf{S}_{W}$ is the single point furthest from $\mathcal{B}_{R}$). At times $t>t_{\sf move}=W/v$, [^43] This is a strict overestimate of the first-move time since e.g. $\mathsf{S}_{0}$ is closer to $\mathcal{B}_{R}$ than to $\mathsf{S}_{W}$, on account of our assumption that $\mathsf{B}_{(L-1-x){\mathbf{\hat{x}}}}(3W/2)\cap\mathcal{B}_{R}\neq\varnothing$. the bias in the feedback rules of Def. 4.3 mean that any anyon in $\mathsf{A}((\mathsf{S}_{k},t))$ must move out of $\mathsf{S}_{k}$ at time $t+1$. If an anyon in $\mathsf{S}_{k}$ at time $t>t_{\sf move}$ moves along ${\mathbf{\hat{x}}}$, then it must move along ${\mathbf{\hat{x}}}$ until annihilating at $\mathcal{B}_{R}$. Therefore by time $t_{\sf move}+W$, every anyon initially in $C$ must have either been annihilated, or begun moving along $+{\mathbf{\hat{x}}}$. Since $x<3W/2$ by assumption, this gives a (crude) upper bound on the erosion time of

### v4

@@ 6.3 Performance guarantees for Lindbladian decoding
$\mathbb{E}[T_{\sf dec}^{\mathsf{a}}(\rho)]=\int_{0}^{\infty}dt\,\mathsf{P}(T_{\sf dec}^{\mathsf{a}}(\rho)>t)\leq\int_{0}^{\infty}dt\,\mathsf{P}(\mathsf{t}_{\sf sim,min}(t)<T_{\sf dec}(\rho)),$ (129)
@@
$\mathbb{E}[T_{\sf dec}^{\mathsf{a}}(\rho)]$ $\leq t_{*}+\int_{t^{*}}^{\infty}dt\,\mathsf{P}(\mathsf{t}_{\sf sim,min}(t)<\gamma_{t}t),$ (130)
@@
$\mathbb{E}[T_{\sf dec}^{\mathsf{a}}(\rho)]$ $=\int_{0}^{\infty}dt\,\mathsf{P}(T_{\sf dec}^{\mathsf{a}}(\rho)>t)\geq\int_{0}^{\infty}dt\,\mathsf{P}(\mathsf{t}_{\sf sim,max}(t)<T_{\sf dec}(\rho))$ (133)
$\geq\int_{a}^{b}dt\,(1-\mathsf{P}(\mathsf{t}_{\sf sim,max}(t)\geq\lambda_{t}t)),$
@@ 7.2 Two dimensions
We now repeat the above analysis for the 2D toric code. The left panel of fig. 11 plots $p_{\sf log}$ as a function of $p$ for different system sizes. Since our simulations are done on the torus, where logical errors occur when anyon strings wrap around either one or both of the two nontrivial cycles, $p_{\sf log}\rightarrow 3/4$ in the limit $p\rightarrow 1/2$. The crossing points for different values of $L$ are consistent with a threshold of $p_{c}\approx 7.3\%$, which unlike the 1D case is (a few percent) smaller than the optimum value of $\approx 11\%$ set by the phase transition of the 2D random bond Ising model on the Nishimori line [21], but is a bit higher than the $\approx 6.1\%$ of the “constant-cc” field-based decoder of Ref. [13]. The right panel of fig. 11 plots $p_{\sf log}$ against $L$ and confirms the expected scaling $p_{\sf log}=e^{-\Theta(L)}$ for $p$ not too far below threshold. Finally, fig. 12 shows $T_{\sf dec}$ as a function of $L$ for several values of $p<p_{c}$, which again display an excellent fit to the predicted scaling $T_{\sf dec}=\Theta(\log L)$ (as in 1D, introducing local spatial correlations into the noise does not appear to affect the results).
@@
We can give an extremely crude estimation of $p_{c}$ for i.i.d noise using the following simplistic argument. Let $\mathsf{P}_{\sf nearest\,\,paired}$ be the probability that a given anyon is paired with one of its nearest neighbors. Heuristically, our decoder should succeed when this probability is >1/2>1/2; in this case, moving each anyon towards its nearest neighbor is likely to induce a pairing leading to an error-free perfect matching. We compute $\mathsf{P}_{\sf nearest\,\,paired}$ numerically in fig. 13, which shows that $\mathsf{P}_{\sf nearest\,\,paired}=1/2$ at $p\approx 6.5\%$, which is slightly smaller than the numerically computed value of the threshold probability. It would be interesting if a refinement of this naive argument could be used to get a sharp bound on $p_{c}$.
@@ 7.3 Uncoordinated desynchronization
For all cellular automata capable of highly structured computations known to the author—such as the game of life and many of the 1D automata defined by Wolfram—performing uncoordinated desynchronization completely destroys the automaton’s ability to perform meaningful computations. However, for Toom’s rule, which like the present decoder is designed only to remember $O(1)$ bits of information, uncoordinated desynchronization is provably innocuous, changing the value of $p_{c}$ but not the existence of a threshold [59, 60, 61]. While it is possible to construct local robust classical memories storing $O(1)$ bits for which uncoordinated desynchronization fails [62], it is still reasonable to conjecture that it may work for the present message-passing decoder under consideration.
@@
Figure 15: Performance of the uncoordinated asynchronous 2D toric code decoder under i.i.d noise of strength pp. Left: $p_{\sf log}^{\sf un}$ vs $p$, consistent with a threshold around $p^{\sf un}_{c}\approx 5.2\%$. Right: $p_{\sf log}^{\sf un}$ vs $p$; the solid lines are fits to $p_{\sf log}=C(p/p^{\sf un}_{c})^{\gamma L}$ with $\gamma\approx 1/10$.
@@ 8 Acknowledgments
I thank Ehud Altman, Nicholas O’Dea, Ruihua Fan, Sarang Gopalakrishnan, Jeongwan Haah, Yaodong Li, Louis Paletta, Pablo Sala, Charles Stahl, Kazuaki Takasan, Zijian Wang, Hayata Yamasaki, and Mike Zaletel for discussions and feedback, Shankar Balasubramanian and Margarita Davydova for helpful discussions about Gacs’ sparsity methods, synchronization, and collaboration on related work, Larry Chen and Noah Goss for insights into hardware constraints, and Nathaniel Selub for finding an egregious number of typos. I am supported by a Miller research fellowship.
@@ A.1 Basic notions, and definitions of level-$k$ noise and error sets
In the following, we will always assume the existence of a finite $0<p<1$ for which the above holds.
@@
It follows from the above definition that any $(W/2,W/2+B)$-isolated point defines a $(W,B)$-cluster. However, a given cluster $C_{(W,B)}$ may contain anywhere between 0 and $|C_{(W,B)}|$ $(W/2,W/2+B)$-isolated points; examples are shown in fig. 16. The partition of a (subset of a) subset of $\mathbb{Z}$ into clusters is as stated not necessarily unique in general, but it is if $B>W$:
@@
Let $\mathsf{N}\subset\mathbb{Z}$ be a noise realization, and $\{C_{(W,B)}^{(i)}\}$ be a collection of $(W,B)$-clusters in $\mathsf{N}$. If $B>W$, these clusters are disjoint: $C_{(W,B)}^{(i)}\cap C_{(W,B)}^{(j)}$ if $i\neq j$.
@@
Fix $i,j$, assume $B>W$, and suppose $x\in C_{(W,B)}^{(i)}\cap C_{(W,B)}^{(j)}$. Consider a point $y\in C_{(W,B)}^{(i)}$. Then |xy|<W|x-y|<W. This implies ${\sf dist}(C_{(W,B)}^{(j)},\{y\})<W$. Thus ${\sf dist}(C_{(W,B)}^{(j)},\mathsf{N}\setminus C_{(W,B)}^{(j)})<B$ since $B>W$, and hence we must have $i=j$. ∎
@@ A.2 The sparsity theorem in one dimension
$\mathsf{E}_{k}=\mathsf{N}_{k}\setminus\mathsf{N}_{k+1}.$ (144)
We similarly define the level-$k$ isolated error set $\mathcal{E}_{k}$ as $\mathcal{E}_{k}=\mathcal{N}_{k}\setminus\mathcal{N}_{k+1}$.
@@
$p_{k}=\max_{x\in\mathbb{Z}}\mathsf{P}\left(x\in\mathsf{N}_{k}\right),$ (147)
@@ A.3 General dimensions
$p_{k}=\max_{\mathbf{x}\in\mathbb{Z}^{D}}\mathsf{P}(\mathbf{x}\in\mathsf{N}_{k}).$ (164)
@@
The sparsity of errors at different values of $k$ is easy enough to verify numerically; the result of doing this is shown in fig. 17. One may also check that the bound on $\mathsf{P}(\mathcal{N}_{k}\neq\varnothing)$ is asymptotically rather tight for strength-$p$ i.i.d noise. As an illustration of this, define $P_{\sf max}$ as the probability to have a system-spanning cluster:
$P_{\sf max}=\mathsf{P}(\mathcal{N}_{\lfloor\log_{n}(L/2\gamma)\rfloor}\neq\varnothing).$ (170)
If the scaling given in the sparsity bound is saturated, we expect
$P_{\sf max}=(p/p_{*})^{\Theta(L^{\log_{n}(2)})}$ (171)
at small enough $p<p_{*}$. This expectation is numerically seen to be satisfied to a reasonable degree; see fig. 18 for details.
@@
Figure 18: $\log(1/P_{\sf max})$ against system size for 1D i.i.d noise of strength pp. The clustering parameters are chosen as $(\beta,\gamma,n)=(3,1,4)$, values for which the sparsity theorem guarantees a threshold, and all three values of $p$ are taken to be sub-threshold. The dashed lines are fits to (171).
@@ Appendix B Open boundary conditions
$\mathcal{B}=\{\mathbf{r}\in\Lambda\,:\,r^{1}=0\}\cup\{\mathbf{r}\in\Lambda\,:\,r^{1}=L-1\}=\mathcal{B}_{L}\cup\mathcal{B}_{R}$ (172)
@@ B.1 Modifications to the decoder
We begin by describing how to modify the message passing and feedback rules in the presence of open boundaries. In this context, error correction is performed by allowing anyons to match either with other anyons, or with sites in $\mathcal{B}$. Since sites in $\mathcal{B}$ act as anyons for the purposes of matching, we simply modify the automaton rule $\mathcal{A}$ to source messages at all such sites, with a component of $\mathcal{B}$ with outward-facing unit normal $^{\mathbf{\hat{a}}}$ sourcing messages that propagate along $\mathbf{\hat{a}}$}. These rules lead to anyons being attracted either to the anyon nearest to them or to the nearest site in $\mathcal{B}$, whichever is closer. This modification is formalized in the following definition:
@@
where we take $C^{(a)}_{\mathbf{r}\mp{\mathbf{\hat{a}}}}$$\varnothing$ if $\mathbf{r}\mp{\mathbf{\hat{a}}}\not\in\Lambda$.
@@ B.2 Erosion with open boundaries
We first claim that if an anyon enters $\Lambda\setminus\mathsf{B}$, it moves along $\mathbf{\hat{x}}$} until annihilating on a site in $\mathcal{B}_{R}$. To show this, consider the first anyon $a$ from $C$ to enter $\Lambda\setminus\mathsf{B}$ (if multiple anyons enter $\Lambda\setminus\mathsf{B}$ at the same time, we may resolve the degeneracy by picking the anyon closest to $\mathcal{B}_{R}$ [or one of them, if there are multiple closest anyons]). If $\mathbf{r}_{t}$ is the spatial coordinate of anyon $a$ at time $t$, then if $t$ is the time when $a$ first moves into $\Lambda\setminus\mathsf{B}$, we must have $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r}_{t},t))=0$ and
[…]
@@
is for there to be an anyon in ${\sf PLF}^{+x}(\mathbf{r}_{t},t)$ that moves along $\mathbf{\hat{x}}$} at time $t$, which contradicts $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r}_{t},t))=0$. Therefore the distance from $a$ to $\mathcal{B}_{R}$ decreases at least as much as the distance between $a$ and any point in $\mathsf{A}({\sf PLF}(\mathbf{r}_{t},t))$. This argument holds for any time $t^{\prime}>t$, and $a$ therefore moves along $\mathbf{\hat{x}}$} until annihilating on a site in $\mathcal{B}_{R}$.
[…]
@@
Therefore $a^{\prime}$ must continue to move along $^{\mathbf{\hat{x}}}$ at the next time step; since this argument holds for all $t^{\prime}$, it shows the claim.
@@
To give a (loose) upper bound on the erosion time, we use the same strategy as in Lemma 4.4. Consider anyons at the slice $\mathsf{S}_{k}=\mathsf{B}\cap\{\mathbf{r}\,:\,r^{1}=L-1-x+W/2-k\}$ (so that $\mathsf{S}_{0}$ is the single point of $\mathsf{B}$ closest to $\mathcal{B}_{R}$, and $\mathsf{S}_{W}$ is the single point furthest from $\mathcal{B}_{R}$). At times $t>t_{\sf move}=W/v$, [^44] This is a strict overestimate of the first-move time since e.g. $\mathsf{S}_{0}$ is closer to $\mathcal{B}_{R}$ than to $\mathsf{S}_{W}$, on account of our assumption that $\mathsf{B}_{(L-1-x){\mathbf{\hat{x}}}}(3W/2)\cap\mathcal{B}_{R}\neq\varnothing$. the bias in the feedback rules of Def. 4.3 mean that any anyon in $\mathsf{A}((\mathsf{S}_{k},t))$ must move out of $\mathsf{S}_{k}$ at time $t+1$. If an anyon in $\mathsf{S}_{k}$ at time $t>t_{\sf move}$ moves along $^{\mathbf{\hat{x}}}$, then it must move along $^{\mathbf{\hat{x}}}$ until annihilating at $\mathcal{B}_{R}$. Therefore by time $t_{\sf move}+W$, every anyon initially in $C$ must have either been annihilated, or begun moving along $\mathbf{\hat{x}}$}. Since $x<3W/2$ by assumption, this gives a (crude) upper bound on the erosion time of